OpenStudy (anonymous):

Anyone that can help me out please? Traffic signs are regulated by the Manual on Uniform Traffic Control Devices (MUTCD). The perimeter of some traffic sign must be 132 inches. Also, its length is 14inches longer than it’s width. Find the dimensions of this sign. The width of the sign is ( ) in. The length of the sign is ( ) in.

OpenStudy (anonymous):

Assume it's a rectangle, then  perimeter is all four sides added together, and length = width +14 inches. Go from there.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Using 'L' for length and 'W' width, tell me in a mathematical way that the perimeter is 132 in.

OpenStudy (anonymous):

\[132=2l + 2w\]\[w+14=l\]\[132=2(w+14)+2w\]\[132=2w+28+2w=4w+28\]\[104=4w\]\[w=26 \]\[l=40\]

OpenStudy (anonymous):

perimeter= L+L+W+W

OpenStudy (anonymous):

thank you all for the help.... and steven, I didn't really want you to blurt out the answer. I wanted to learn....but thank you too.

OpenStudy (anonymous):

@smstevns http://openstudy.com/code-of-conduct